the radius of the field in the form of a sector is 21 meter .thecost of constructing of wall around hte field is 1875 rs. at the rate of 25 per meter if is cost 10 per m^2 to till the field ,what will be the cost of tilling the whole filed?
1
Expert's answer
2013-02-27T06:56:54-0500
Perimeter of the sector is r+r+r*a, where a is angle of the sector. Then25(2r+ra)=1875, r=21, 21(2+a)=75, and we have a=75/21 - 2 = 25/7 - 2 = 11/7
Area S=a*r*r/2 = 441 * 11/14=346.5.